# uGMM-NN: Univariate Gaussian Mixture Model Neural Network

### v1

@@ Abstract
This paper introduces the Univariate Gaussian Mixture Model Neural Network (uGMM-NN), a novel neural architecture that embeds probabilistic reasoning directly into the computational units of deep networks. Unlike traditional neurons, which apply weighted sums followed by fixed nonlinearities, each uGMM-NN node parameterizes its activations as a univariate Gaussian mixture, with learnable means, variances, and mixing coefficients. This design enables richer representations by capturing multimodality and uncertainty at the level of individual neurons, while retaining the scalability of standard feedforward networks. We demonstrate that uGMM-NN can achieve competitive discriminative performance compared to conventional multilayer perceptrons, while additionally offering a probabilistic interpretation of activations. The proposed framework provides a foundation for integrating uncertainty-aware components into modern neural architectures, opening new directions for both discriminative and generative modeling.
Keywords: probabilistic neurons, deep learning, probabilistic circuits, Gaussian mixture, uGMM, uGMM-NN
@@ 1 Introduction
In contrast, this paper introduces a novel approach by embedding a univariate Gaussian Mixture Model (uGMM) directly into the network’s computational units, enabling each neuron to represent a multimodal univariate Gaussian distribution, thereby directly addressing the limitations of deterministic neurons from a new perspective. While our primary focus is on discriminative applications where the goal is to predict a class label, the probabilistic backbone of our model shares conceptual similarities with probabilistic circuits. This suggests that when trained generatively, the model could potentially learn complex joint distributions. We evaluate uGMM-NN on two representative datasets. On the image dataset MNIST, we compare its discriminative performance against a standard Multi-Layer Perceptron (MLP). On the tabular dataset, Iris, we highlight the model’s ability to capture joint distributions for inference, drawing conceptual connections to probabilistic graphical models and probabilistic circuits.
@@ 3.1 Univariate GMM Nodes
Let $\mathbf{x}=[x_{1},\dots,x_{N}]$ denote the inputs to neuron $j$ from the previous layer, where $N$ is the number of neurons in that layer. In uGMM-NN, each neuron $j$ defines a univariate Gaussian mixture over a latent variable $y$, parameterized by learnable means $\mu_{j,k}$, variances $\sigma^{2}_{j,k}$, and mixing coefficients $\pi_{j,k}$ for each input $x_{k}$:
$P_{j}(y)=\sum_{k=1}^{N}\pi_{j,k}\,\mathcal{N}(y\mid\mu_{j,k},\sigma^{2}_{j,k}),\quad\text{with}\quad\sum_{k=1}^{N}\pi_{j,k}=1.$
Here, each input contributes a component to the mixture distribution. The number of components $N$ is thus determined by the width of the previous layer, not treated as an independent hyperparameter.
@@
$P_{j}(y)=\sum_{k=1}^{N}\pi_{j,k}\cdot\frac{1}{\sqrt{2\pi\sigma^{2}_{j,k}}}\exp\left(-\frac{(y-\mu_{j,k})^{2}}{2\sigma^{2}_{j,k}}\right).$
@@
$\log P_{j}(y)=\log\left(\sum_{k=1}^{N}\exp\left[\log\pi_{j,k}-\frac{1}{2}\log(2\pi\sigma^{2}_{j,k})-\frac{(y-\mu_{j,k})^{2}}{2\sigma^{2}_{j,k}}\right]\right),$
@@
Figure 1: Schematic of a uGMM-NN neuron. Each neuron models a uGMM with NN components based on inputs from the previous layer, with each connection parameterized by mean ($\mu_{j,k}$), standard deviation ($\sigma_{j,k}$), and mixing coefficient ($\pi_{j,k}$), outputting the log probability $\log P_{j}(y)$.
@@ 3.2 Probabilistic Interpretability of Neurons
Figure 2 illustrates a single uGMM neuron with three mixture components. Each component is parameterized by a mean $\mu_{j,k}$, variance $\sigma_{j,k}^{2}$, and mixing coefficient $\pi_{j,k}$. Together, these define the neuron’s overall probability density $P_{j}(y)$, with the dashed black line representing the full mixture. The neuron’s output is the log-density $\log P_{j}(y)$, which we interpret as its activation.
This formulation enables an intuitive understanding of what the neuron ”detects.” For example, a component with a high mixing coefficient $\pi_{j,k}$ indicates that the corresponding input $x_{k}$ plays a dominant role in shaping the neuron’s density. Similarly, the spread of the Gaussian (via $\sigma_{j,k}^{2}$) reflects how confidently the neuron attributes input mass to a given region of its input space. From this perspective, each neuron defines a probabilistic partition over its input domain, where mixture components correspond to interpretable submodes of the neuron’s response. This is in contrast to traditional activations, which typically lack such localized semantics. As a result, uGMM-NN offers built-in interpretability, where the internal representation of each layer is not just a scalar value but a structured probability landscape.
@@
Figure 2: Illustration of a single uGMM neuron with three univariate Gaussian mixture components. The means ($\mu_{j,k}$), variances ($\sigma_{j,k}^{2}$), and mixing coefficients ($\pi_{j,k}$) define the neuron’s probability density $P_{j}(y)$. The dashed black line shows the combined mixture density.
@@ 3.3 Network Structure and Training
The uGMM-NN is organized in a standard feedforward manner with input, hidden, and output layers. Each neuron in a hidden layer computes a univariate Gaussian mixture over the activations of the previous layer, producing a log-density that serves as input for the next layer. Importantly, the number of mixture components in each neuron is automatically determined by the width of the preceding layer, ensuring a direct mapping from inputs to probabilistic components.
Stacking multiple uGMM layers allows the network to build hierarchical probabilistic representations, capturing increasingly complex and multimodal patterns in the data. Unlike traditional MLPs, where each neuron performs a 1D affine transformation followed by a nonlinearity and outputs a deterministic scalar, uGMM neurons operate on univariate inputs and output log-probability values under a learned Gaussian mixture. These can be viewed as 1D probabilistic operations, analogous in shape and dimension to their MLP counterparts. By focusing on univariate GMMs at each neuron, uGMM-NN avoids the exponential complexity often associated with multivariate mixture models, allowing it to scale efficiently to deep architectures. This architectural simplicity enables practical training and inference on large datasets while retaining the expressive power of probabilistic modeling.
@@ 3.4 Dropout in uGMM-NN
where $p$ is the dropout probability. Components that are ”dropped” are effectively removed from the mixture by setting their log-density contribution to -$\infty$. The resulting log-density with dropout is
$\log P_{j}(y)_{\text{dropout}}=\log\sum_{k=1}^{N}\pi_{j,k}\,d_{j,k}\,\mathcal{N}(y\mid\mu_{j,k},\sigma_{j,k}^{2}),$
where $\pi_{j,k}$, $\mu_{j,k}$, and $\sigma_{j,k}^{2}$ are the mixture coefficients, means, and variances of the $k$-th component of neuron $j$, respectively. Dropout is applied only during training and is disabled during inference, allowing the full mixture to contribute to the neuron’s output.
@@ 3.5 Universal Approximation Perspective
In comparison, standard MLPs and Kolmogorov–Arnold Networks (KANs) provide universal approximation guarantees for deterministic functions. uGMM-NN extends this notion to probabilistic, multimodal representations, combining expressive power with interpretable uncertainty through mixture parameters $(\pi_{j,k},\mu_{j,k},\sigma_{j,k}^{2})$.
@@ 4 Experiments
We empirically evaluate the uGMM-NN on classification benchmarks, comparing it against a standard Feedforward Neural Network (FFNN) with equivalent architecture. The FFNN consists of linear layers with ReLU activations, while the uGMM-NN replaces each neuron with a uGMM node as described in Section 3.1. To ensure a fair comparison, both models were matched in hidden layer sizes and total depth. Dropout was applied in all models to reduce overfitting, and hyperparameters were selected based on validation performance.
@@ 4.1 Datasets
We considered two widely used benchmarks:
  • MNIST [14]: A dataset of 70,000 grayscale images (60,000 train, 10,000 test) of handwritten digits ($\times28$ pixels), with 10 output classes.
  • Iris [15]: A classic dataset of 150 samples with 4 numerical features and 3 output classes, corresponding to different iris species.
@@ 4.2 Training Setup
All models were trained using the Adam optimizer with default parameters ($\beta_{1}=0.9,\beta_{2}=0.999$), and an initial learning rate tuned in $\{10^{-4},10^{-3},10^{-2}\}$. Batch normalization was applied in the FFNN baselines where applicable.
Both models were trained for $100$ epochs with the following learning rate schedule:
$\texttt{MultiStepLR}(\texttt{milestones}=[20,45,60],\ \gamma=0.1).$ (1)
The FFNN used an initial learning rate of $10^{-3}$ and applied dropout with probability $p=0.5$ on the 64-unit layer. The uGMM-NN used a larger initial learning rate of $10^{-2}$, reflecting the faster adaptation required by mixture parameters, and applied dropout with probability $p=0.3$ on the 128-unit layer.
For uGMM-NN, the mixture parameters $(\mu_{j,k},\sigma_{j,k}^{2},\pi_{j,k})$ were initialized randomly, and gradients were backpropagated through the log-density computations using automatic differentiation. On the Iris dataset, uGMM-NN was trained generatively to maximize the joint likelihood of features and labels, with predictions obtained by posterior inference. On MNIST, both uGMM-NN and FFNN were trained discriminatively using cross-entropy loss at the output layer.
@@ 4.3 Performance
Results are reported in Table 1.
On the Iris dataset, uGMM-NN was trained generatively to maximize the joint likelihood of features and labels. Predictions were obtained via posterior inference:
@@
where $\mathbf{x}$ is the feature vector. The FFNN baseline was trained discriminatively with cross-entropy loss. Both models achieved perfect classification accuracy, illustrating that uGMM-NN can perform competitively on tabular datasets even when trained generatively. While the Iris dataset is relatively simple, these results motivate future evaluation of uGMM-NN on more complex tabular benchmarks.
On MNIST, both models converged to strong performance after 100 epochs. The FFNN achieved a stable test accuracy of $98.21\%$ (error rate $1.79\%$) in the final training window (epochs 90–100). The uGMM-NN attained a slightly lower but competitive test accuracy of $97.74\%$ (error rate $2.26\%$) across the same interval. This indicates that while the FFNN maintains a small advantage in raw discriminative accuracy, the uGMM-NN remains within $0.5\%$ of the baseline while offering additional benefits such as uncertainty quantification and a probabilistic interpretation of hidden activations.
Table 1: Test accuracy and training loss for uGMM-NN vs. FFNN across benchmark datasets.
Dataset Model Test Accuracy (%) Training Loss
Iris FFNN 100.0 $\pm$ 0.0 Cross-Entropy
Iris uGMM-NN 100.0 $\pm$ 0.0 NLL (Generative)
MNIST FFNN 98.21 Cross-Entropy
MNIST uGMM-NN 97.74 Cross-Entropy

### v2

@@ Abstract
This paper introduces the Univariate Gaussian Mixture Model Neural Network (uGMM-NN), a novel architecture that embeds probabilistic reasoning directly into deep networks. Unlike traditional neurons, which apply weighted sums followed by fixed nonlinearities, each node within the uGMM-NN is parameterized by a univariate Gaussian mixture, with learnable means, variances, and mixing coefficients. This design enables richer representations by capturing multimodality and uncertainty at the level of individual neurons, while retaining the scalability of standard feedforward networks. We demonstrate that uGMM-NN can achieve competitive discriminative performance compared to conventional multilayer perceptrons, while additionally offering a probabilistic interpretation of activations. The proposed framework provides a foundation for integrating uncertainty-aware components into modern neural architectures, opening new directions for both discriminative and generative modeling.
Keywords: probabilistic neurons, deep learning, probabilistic circuits, Gaussian mixture, uGMM Layer, uGMM, uGMM-NN
@@ 1 Introduction
In contrast, this paper introduces a novel approach by embedding a univariate Gaussian Mixture Model (uGMM) directly into the network’s computational units, enabling each neuron to represent a multimodal univariate Gaussian distribution, thereby directly addressing the limitations of deterministic neurons from a new perspective. While our primary focus is on discriminative applications where the goal is to predict a class label, the probabilistic backbone of our model shares conceptual similarities with probabilistic circuits. We evaluate the uGMM-NN across two representative datasets and two architectural paradigms:
  1. 1.
    On MNIST, we perform a direct comparison of discriminative performance between the uGMM-NN and a standard Multi-Layer Perceptron (MLP), and further demonstrate the versatility of our approach by replacing the fully connected layers of a convolutional network with uGMM neurons (CNN-uGMM vs. CNN-MLP).
  2. 2.
    On the classic Iris dataset, we offer a preliminary demonstration of the model’s capacity for generative modeling of a joint distribution. This experiment serves to illustrate the conceptual connection between uGMM-NN and probabilistic graphical models.
@@ 2 Related Work
Radial Basis Function (RBF) networks also employ Gaussian-shaped responses at the unit level [11], but they are not probabilistic models: their activations are unnormalized basis functions centered at fixed prototypes, without mixture weights that sum to one or a likelihood-based interpretation. By contrast, each uGMM neuron defines a normalized Gaussian mixture with trainable means, variances, and mixing weights, and outputs a log-density. Consequently, despite a superficial similarity in functional form, uGMM-NN is conceptually closer to probabilistic circuits than to classical RBF networks.
@@ 3.1 Univariate GMM Nodes
Let $\mathbf{x}=[x_{1},\dots,x_{N}]$ denote the inputs to neuron $j$ from the previous layer, where $N$ is the number of neurons in that layer. In uGMM-NN, each neuron $j$ defines a univariate Gaussian mixture over a latent variable $y$, parameterized by learnable means $\mu_{j,k}$, variances $\sigma^{2}_{j,k}$, and mixing weights $w_{j,k}$ for each input $x_{k}$:
$P_{j}(y)=\sum_{k=1}^{N}w_{j,k}\,\mathcal{N}(y\mid\mu_{j,k},\sigma^{2}_{j,k}),\quad\text{with}\quad\sum_{k=1}^{N}w_{j,k}=1.$
Each input $x_{k}$ contributes one component to the mixture. Therefore, the number of mixture components $N$ is determined by the width of the previous layer, rather than chosen as a separate hyperparameter.
@@
$P_{j}(y)=\sum_{k=1}^{N}w_{j,k}\cdot\frac{1}{\sqrt{2\pi\sigma^{2}_{j,k}}}\exp\left(-\frac{(y-\mu_{j,k})^{2}}{2\sigma^{2}_{j,k}}\right).$
@@
$\log P_{j}(y)=\log\left(\sum_{k=1}^{N}\exp\left[\log w_{j,k}-\frac{1}{2}\log(2\pi\sigma^{2}_{j,k})-\frac{(y-\mu_{j,k})^{2}}{2\sigma^{2}_{j,k}}\right]\right),$
@@
Figure 1: Schematic of a uGMM-NN neuron. Each neuron models a uGMM with NN components based on inputs from the previous layer, with each connection parameterized by mean ($\mu_{j,k}$), standard deviation ($\sigma_{j,k}$), and mixing weight ($w_{j,k}$), outputting the log probability $\log P_{j}(y)$.
@@ 3.2 Probabilistic Interpretability of Neurons
Figure 2 illustrates a single uGMM neuron with three mixture components. Each component is parameterized by a mean $\mu_{j,k}$, standard deviation $\sigma_{j,k}$, and mixing weight $w_{j,k}$ (with the variance $\sigma_{j,k}^{2}$ used in the log-density computation). Together, these define the neuron’s overall probability density $P_{j}(y)$, with the dashed black line representing the full mixture. The neuron’s output is the log-density $\log P_{j}(y)$, which we interpret as its activation.
This formulation enables an intuitive understanding of what the neuron ”detects.” For example, a component with a high mixing weight $w_{j,k}$ indicates that the neuron assigns significant probability mass to the corresponding region of its input space. Similarly, the spread of the Gaussian (via $\sigma_{j,k}^{2}$) reflects how confidently the neuron attributes input mass to a given region of its input space. From this perspective, each neuron defines a probabilistic partition over its input domain, where mixture components correspond to interpretable submodes of the neuron’s response. This is in contrast to traditional activations, which typically lack such localized semantics. As a result, uGMM-NN offers built-in interpretability, where the internal representation of each layer is not just a scalar value but a structured probability landscape.
@@
Figure 2: Illustration of a single uGMM neuron with three univariate Gaussian mixture components. The means ($\mu_{j,k}$), variances ($\sigma_{j,k}^{2}$), and mixing weights ($w_{j,k}$) define the neuron’s probability density $P_{j}(y)$. The dashed black line shows the combined mixture density.
@@ 3.3 Network Structure and Training
The uGMM-NN is organized in a standard feedforward manner with input, hidden, and output layers. Each hidden layer can be interpreted as a Gaussian mixture layer, where every neuron computes a univariate Gaussian mixture model over the activations of the previous layer, producing log-densities that serve as inputs for the next layer. Importantly, the number of mixture components in each neuron is automatically determined by the width of the preceding layer, ensuring a direct mapping from inputs to probabilistic components.
Stacking multiple univariate Gaussian Mixture Model (uGMM) layers allows the network to build hierarchical probabilistic representations, capturing increasingly complex and multimodal patterns in the data. Unlike traditional MLPs, where each neuron performs a 1D affine transformation followed by a nonlinearity and outputs a deterministic scalar, uGMM neurons operate on univariate inputs and output log-probability values under a learned Gaussian mixture. These can be viewed as 1D probabilistic operations, analogous in shape and dimension to their MLP counterparts. By focusing on univariate GMMs at each neuron, uGMM-NN avoids the exponential complexity often associated with multivariate mixture models, allowing it to scale efficiently to deep architectures. This architectural simplicity enables practical training and inference on large datasets while retaining the expressive power of probabilistic modeling.
@@ 3.4 Dropout in uGMM-NN
where $p$ is the dropout probability. Components that are ”dropped” are effectively removed from the mixture by setting their log-density contribution to -$\infty$. The resulting log-density with dropout is:
$\log P_{j}(y)_{\text{dropout}}=\log\sum_{k=1}^{N}w_{j,k}\,d_{j,k}\,\mathcal{N}(y\mid\mu_{j,k},\sigma_{j,k}^{2}),$
where $w_{j,k}$, $\mu_{j,k}$, and $\sigma_{j,k}^{2}$ are the mixture weights, means, and variances of the $k$-th component of neuron $j$, respectively. Dropout is applied only during training and is disabled during inference, allowing the full mixture to contribute to the neuron’s output.
@@ 3.5 Universal Approximation Perspective
In comparison, standard MLPs and Kolmogorov–Arnold Networks (KANs) provide universal approximation guarantees for deterministic functions. uGMM-NN extends this notion to probabilistic, multimodal representations, combining expressive power with interpretable uncertainty through mixture parameters $(w_{j,k},\mu_{j,k},\sigma_{j,k}^{2})$.
@@ 4 Experiments
We conduct a series of experiments to evaluate the performance and properties of uGMM-NN, focusing on both generative and discriminative tasks, and comparing it to standard MLP architectures. Specifically, we perform the following experiments:
  1. 1.
    Generative Learning on Iris: A proof-of-concept experiment where the uGMM-NN is trained to model the joint probability distribution of the features and labels, demonstrating its ability to capture multimodal distributions and perform simple posterior inference.
  2. 2.
    Discriminative Comparison on MNIST: A direct comparison between a standard fully-connected MLP and a uGMM-NN of equivalent depth and hidden size, evaluating classification accuracy and training stability on a classic image classification benchmark.
  3. 3.
    Convolutional Architectures: Extending the comparison to convolutional networks, we replace the dense layers in a CNN with uGMM neurons (CNN-uGMM) and compare performance against a conventional CNN with standard MLP layers (CNN-MLP). This demonstrates the versatility of uGMM neurons in modern architectures.
In the MNIST classification experiments, dropout is applied consistently to prevent overfitting, and hyperparameters are selected based on validation performance to ensure fair comparison between models.
@@ 4.1 Datasets
We considered two datasets for our evaluation:
  • MNIST [15]: A common benchmark for deep learning image classification, consisting of 70,000 grayscale images (60,000 train, 10,000 test) of handwritten digits ($\times28$ pixels), with 10 output classes.
  • Iris [16]: A classic small-scale dataset often employed for foundational classification tasks, comprising 150 samples with 4 numerical features and 3 output classes, corresponding to different iris species.
@@ 4.2 Training Setup
This section details the training methodologies and hyperparameters used for our models.
4.2.1 Generative Training
For the generative experiment on the Iris dataset, the network was trained to model the joint probability distribution of the features and labels, $P(\mathbf{x},y)$. The architecture consisted of two simple hidden layers with 10 and 5 uGMM neurons, respectively. The network outputs a single log-density value, representing the log-likelihood of the input.
The model was trained to maximise the average log-likelihood of the training data, using the Adam optimizer with a learning rate of $10^{-3}$. No dropout was applied, and training proceeded for 300 epochs to ensure convergence.
4.2.2 Discriminative Training
In the MNIST classification experiments, all models were trained using the Adam optimizer with default parameters ($\beta_{1}=0.9,\beta_{2}=0.999$) and a cross-entropy loss function. Training was performed for 100 epochs using the MultiStepLR learning rate scheduler with a decay factor of $\gamma=0.1$.
For the fully connected setting, we compared a standard MLP with a uGMM-NN of equivalent depth and hidden layer sizes. For the uGMM-NN, mixture parameters $(\mu_{j,k},\sigma_{j,k}^{2},w_{j,k})$ were initialized randomly, and dropout was applied consistently to prevent overfitting. Hyperparameters, including dropout probability and learning rate schedule milestones, were selected based on validation performance to ensure a fair comparison between models.
Similarly, for the convolutional models, we trained a conventional CNN with MLP layers (CNN-MLP) and a CNN where the dense layers were replaced with uGMM neurons (CNN-uGMM). Dropout and other hyperparameters were chosen analogously to the fully-connected setting to maintain a consistent training protocol across architectures. Table 1 summarizes the optimized hyperparameters for all MNIST discriminative models.
Table 1: Optimized Training hyperparameters for MNIST Discriminative Models
Model Initial Learning Rate Dropout $p$ (Layer) LR Milestones
MLP $10^{-3}$ $0.5$ (128-unit) $[20,45]$
uGMM-NN $10^{-2}$ $0.5$ (128-unit) $\mathbf{[40]}$
CNN-MLP $10^{-3}$ $0.5$ (128-unit) $[20,45]$
CNN-uGMM $10^{-3}$ $0.5$ (128-unit) $\mathbf{[45,80]}$
The uGMM-NN utilised a larger initial learning rate and adjusted learning-rate milestones, reflecting the faster adaptation required by its mixture parameters. Moreover, a consistent dropout rate of $p=0.5$ was applied across all models to ensure comparable regularisation.
@@ 4.3 Performance
When trained generatively, the uGMM-NN model achieved perfect classification accuracy on the Iris dataset. Predictions were obtained via posterior inference:
@@
where $\mathbf{x}$ is the feature vector. This strong performance on a simple, foundational dataset demonstrates its capability to perform discriminative inference from a generatively trained probabilistic model. These results motivate future evaluation of uGMM-NN on more complex tabular benchmarks, though the primary focus of this paper remains on its discriminative abilities, particularly in comparison to traditional MLPs.
For the fully connected models on MNIST, both the MLP and the uGMM-NN converged to strong performance after 100 epochs. The MLP achieved a stable test accuracy of $98.15\%$ over the final training window (epochs 90–100), as reported in Table 2. The uGMM-NN initially reached a slightly lower but still competitive accuracy of $97.8\%$. This small gap is consistent with the wider dynamic range and higher variance of the log-density activations produced by uGMM neurons, which propagate richer statistical information than the deterministic scalar activations of the MLP. To explicitly address this effect and stabilise activation scales, Layer Normalization  [17] was applied to the uGMM-NN. Although Layer Normalization slowed the initial learning trajectory, it ultimately enabled the model to reach a higher and more stable accuracy of $98.34\%$. These results demonstrate that, when the statistical characteristics of its probabilistic activations are properly managed, the uGMM-NN can match or marginally exceed the MLP’s discriminative performance while additionally offering uncertainty quantification and improved interpretability.
For the convolutional models on MNIST, all architectures achieved strong performance, with test accuracies approaching or exceeding $99\%$. The baseline CNN-MLP reached a test accuracy of $99.35\%$, reflecting the well-established effectiveness of convolutional architectures on this dataset.
Replacing the final fully connected layers with uGMM neurons resulted in a CNN-uGMM model that achieved an accuracy of $98.77\%$. When the same normalization strategy described above was applied, the CNN-uGMM reached a higher accuracy of $99.12\%$, demonstrating that uGMM layers can be integrated into convolutional architectures while preserving competitive discriminative performance. Importantly, this substitution introduces a transparent, probabilistic, and structurally interpretable alternative to the standard dense layer.
Figure 3(b) compares the test loss convergence of the CNN-uGMM and CNN-MLP architectures over 100 training epochs. Although final accuracy is comparable, the optimisation dynamics differ markedly. The deterministic CNN-MLP shows the expected smooth and rapid decline in cross-entropy loss, reaching a stable low value within the first 10 epochs. In contrast, the CNN-uGMM variants exhibit a steeper and more variable early trajectory, with substantially higher initial loss and pronounced oscillations during the first 20 epochs. This behaviour reflects the added stochasticity introduced by the uGMM log-density activations, which propagate distributional information rather than fixed point activations.
(a) Fully-connected (uGMM-NN vs MLP) training on MNIST.
(b) Convolutional (CNN-uGMM vs CNN-MLP) training on MNIST.
Figure 3: Comparison of test loss convergence between MLP, uGMM-NN, and their LayerNorm-enhanced or CNN-augmented counterparts. The uGMM layers achieve comparable or superior final loss and stability, particularly when Layer Normalization is applied to stabilize the log-density activations.
Table 2: Test accuracy and training loss for uGMM-NN vs. MLP across benchmark datasets.
Dataset Model Test Accuracy (%)
MNIST MLP 98.15
MNIST uGMM-NN 97.8
MNIST uGMM-NN + LayerNorm 98.34
MNIST CNN-uGMM 98.77
MNIST CNN-uGMM + LayerNorm 99.12
MNIST CNN-MLP 99.35